Exercise 11-4 Direct Labor and Variable Manufacturing Overhead Variances [LO11-2, LO11-3] Erie Company manufactures a small mp3 player called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate mp3 player are as follows: Standard Hours Standard Rate per Hour Standard Cost 30 minutes $5.80 $2.90 During August, 10,630 hours of direct labor time were needed to make 19,800 units of the Jogging Mate. The direct labor cost totaled $59,528 for the month. Required: 1. According to the standards, what direct labor cost should have been incurred to make 19,800 units of the Jogging Mate? By how much does this differ from the cost that was incurred? (Round Standard labor time per unit and Standard direct labor rate to 2 decimal places.) 2. Break down the difference in cost from (1) above into a labor rate variance and a labor efficiency variance. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations.) 3. The budgeted variable manufacturing overhead rate is $4.5 per direct labor-hour. During August, the company incurred $51,024 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations and round your final answers to nearest whole dollar.)
Answer #1

1.

Total standard cost = Standard cost per unit X Number of units = 2.90 X 19,800 = $ 57,420

Actual costs incurred = $ 59,528

Variance = 59,528 - 57,420 = $ 2,108 UNFAVORABLE

2 and 3

Standard Actual
Qty Rate Amount Qty Rate Amount
Labor 9,900 5.80 57,420 Labor 10,630 5.60 59,528
Overhead 9,900 4.5 44,550 Overhead 10,630 4.80 51,024
Labor Overhead
Efficiency variance = (Actual Qty - Standard Qty) X Standard Rate =(10,630- 9,900) X 5.8 $ 4,234 Unfavorable =(10,630- 9,900) X 4.5 $ 3,285 Unfavorable
Rate Variance = (Actual Rate - Standard Rate) X Actual Qty =(5.60 - 5.80) X 10,630 $(2,126) Favorable =(4.80 - 4.50) X 10,630 $ 3,189 Unfavorable
$ 2,108 Unfavorable $ 6,474 Unfavorable